OpenStudy (anonymous):

i need help asap Identify each expression as equivalent to either 2(3x + 7y) or 1/2(12x + 14y).

OpenStudy (anonymous):

@Destinymasha

OpenStudy (anonymous):

@sircange

OpenStudy (mathstudent55):

You need to know if the two expressions above are equivalent?

OpenStudy (mathstudent55):

To find out, distribute the 2 by the left one, and distribute the 1/2 by the right one.

OpenStudy (anonymous):

so whats the answer?

OpenStudy (mathstudent55):

Do you know how to use the distributive property?

OpenStudy (anonymous):

no

OpenStudy (mathstudent55):

Multiply the number outside the parentheses by each term inside the parentheses. For example: 3(2x + 4) = 3 * 2x + 3 * 4 = 6x + 12

OpenStudy (mathstudent55):

Now let's do it with your expressions: 2(3x + 7y) 1/2(12x + 14y) =2 * 3x + 2 * 7y =1/2 * 12x + 1/2 * 14y = 6x + 14y = 6x + 7y

OpenStudy (mathstudent55):

Now you can answer whether the expressions are equivalent or not.

OpenStudy (anonymous):

so no because the seacond one is 6x +7y and the first one is 6x + 14y

OpenStudy (mathstudent55):

Right. They are not equivalent.

OpenStudy (anonymous):

thanks i have a couple more can you help

OpenStudy (mathstudent55):

Sorry, I need to go in about 5 minutes. If you have a quick one, I'll try. Pls start a new post.